OpenStudy (anonymous):

At a sand plant, sand is falling off a conveyor and onto a conical pile at the rate of 10 cubic feet per minute. The diameter of the base of the cone is approximately three times the altitude. At what rate is the height of the pile changing when it is 15 feet high? Answer: .006 ft/min (8/405pi)

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Here are a lot of examples of this type problem http://tutorial.math.lamar.edu/Classes/CalcI/RelatedRates.aspx

OpenStudy (anonymous):

oh ok thx but can u show me how to do this one particular problem? so far i put d=3h 2r=3h \[r=\frac{ 3 }{ 2 }h\]

OpenStudy (anonymous):

my internet is kinda laggy :x

OpenStudy (phi):

Start with the volume of a cone \[ V= \frac{1}{3} \pi r^2 h\] they say the diameter is 3 times the height

OpenStudy (phi):

they give dV change in volume they want dh change in height. so we want to replace r with h in the formula

OpenStudy (anonymous):

so we take the derivative of \[V=\frac{ 1 }{ 3}\pi r^2h\] in terms of h? :)

OpenStudy (phi):

first plug in r= 3/2 h, then take the derivative

OpenStudy (anonymous):

after i plug in r i take the derivative in terms of "h"?

OpenStudy (phi):

\[ V=\frac{ 1 }{ 3}\pi r^2h \] with r= 3/2 h you get \[ V=\frac{ 3 }{ 4}\pi h^3 \] now take the derivative \[ dV=\frac{ 3 }{ 4}\pi d\left(h^3\right) \]

OpenStudy (phi):

better, take the derivative with respect to time \[ \frac{dV}{dt}=\frac{ 3 }{ 4}\pi \frac{d}{dt}\left(h^3\right)\]

OpenStudy (anonymous):

ok then what

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

how did u get "d" in \[dV = \frac{ 3 }{ 4 }\pi d{}(h^3)\]

OpenStudy (anonymous):

where did the random d come from

OpenStudy (phi):

I meant to say derivative wrt time

OpenStudy (anonymous):

how do i get h by itself now

OpenStudy (phi):

did you take the derivative? what did you get?

OpenStudy (anonymous):

i take the derivative of \[\frac{ 3 }{ 4 }\pi h^3\] rite?

OpenStudy (phi):

take the derivative of both sides of the equation with respect to time

OpenStudy (anonymous):

how do i do that

OpenStudy (phi):

\[ \frac{dV}{dt}=\frac{ 3 }{ 4}\pi \frac{d}{dt}\left(h^3\right) \] d/dt of V is dV/dt d/dt (h^3) is 3 h^2 dh/dt

OpenStudy (anonymous):

like this? \[\frac{ dV }{ dt } = \frac{ 3 }{ 4 }\pi 3h^2\]

OpenStudy (anonymous):

@phi

OpenStudy (phi):

with these problems you cannot leave out the dh/dt part if you take the derivative of x^3 with respect to x you do 3 x^2 dx/dx and the dx/dx is 1. you get 3x^2 but if you take the derivative with respect to t you do 3 x^2 dx/dt and the dx/dt stays

OpenStudy (anonymous):

so how do i get h by itself

OpenStudy (phi):

you don't. you should now have \[ \frac{dV}{dt}=\frac{ 9 }{ 4}\pi h^2 \frac{dh}{dt}\] now you plug in what you know and solve for dh/dt, the rate that the height is changing

OpenStudy (anonymous):

answer was .006ft/min

OpenStudy (phi):

dv/dt = 10 cu ft/min h= 15 ft dh/dt will be in ft/min

OpenStudy (phi):

\[ \frac{dh}{dt}= 10 \frac{4}{\pi \cdot 9 \cdot 15^2} \]

OpenStudy (phi):

it's algebra at this point.

OpenStudy (anonymous):

how did u get \[\frac{ dh }{ dt }=10\frac{ 4 }{ \pi \times 9 \times 15^2 }\]

OpenStudy (phi):

look at my post where we have the derivative dV/dt = ... we know dv/dt, and h. sub in their values, and solve for dh/dt

OpenStudy (anonymous):

ohh okay ty :)

OpenStudy (anonymous):

can u help me with another problem? ^^

OpenStudy (phi):

ok, please post it.